OpenStudy (howard-wolowitz):

@StudyGurl14 @jdoe0001

OpenStudy (howard-wolowitz):

Ok, could I do this one the same way?

OpenStudy (jdoe0001):

\(\textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} \sqrt{ x + {\color{red}{ 2}} } + {\color{green}{ D}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array}\)

OpenStudy (jdoe0001):

should be simple enough

OpenStudy (howard-wolowitz):

would it switch to negative

OpenStudy (welshfella):

f(x) ---> f(x+a) is a transformation of a units to the left

OpenStudy (jdoe0001):

welll. just notice if Cis negative or positive and take it from there :)

OpenStudy (howard-wolowitz):

ok I got D

OpenStudy (studygurl14):

Yep.

OpenStudy (howard-wolowitz):

plz tell me thats right

OpenStudy (jdoe0001):

yeap, +2 means 2 units to the left